OpenStudy (anonymous):

A number decreased by negative three is equal to -40. Write an equation and find the number.

OpenStudy (texaschic101):

x -(-3) = -40 x + 3 = -40 x = -40 + 3 x = -37 you better get a second opinion on this....I am not sure

OpenStudy (anonymous):

x-(-3) = -40.

OpenStudy (texaschic101):

thats what I thought.....thanks for the confirmation